Solve for:
6x + x(x - 4)
Instructions: Simplify, type exact answers using radicals as needed, express numbers in terms of i and use a comma to separate answers.
Thank you!!

Answer by checkley77(12844)

A QUADRATIC EQUATION REQUIRES AN (=) SIGN.
6x + x(x - 4)=0
6X+X^2-4X=0
X^2-4X+6X=0
X^2+2X=0
X(X+2)=0
X=0 ANS.
X+2=0
X=2 ANS.